17-18 October 2008

Observers: DThompson, JHill, ARakich in Tucson, LUCIFER folks (N.Ageorges,W.Seifert, et al)
Software Support: PGrenz, TEdgin in Tucson
Telescope Operators: ACeranski

Summary

Details

Initial startup LUCIFER working on SW issues.

Update Collimation and Pointing Models

03:20 UT : JMH has updated the LUCIFER/IRTC collimation models to reflect: a) the relative secondary and primary positions that Andrew and Doug have been using the last 4 nights while measuring the field aberrations, and b) the initial collimation correction to remove coma (apparently the 20081007 model had a thermal gradient on the structure).

The new global offsets for M1 are X=-0.8 Y=-2.00 Z=+0.82 RX=-54 RY=-2 in SXPMLUCIFERCollimation.dat. Note that because the active correction has been included, this has reversed the sign of the Y-value compared to what we have been using.

The new global offsets for M2 are X=-7.01 Y=-1.70 Z=0 RX=222.3 RY-181.8 in SXSMLUCIFERCollimation.dat.

The above changes introduce a 40 arcsec change in elevation pointing, so the pointing model has been adjusted to IE=-14.91 CA=-10.65 in LUCIFERSX.20081017.ptmod.

LUCIFER Offset Tests

UT03:43 Problem with spurious astigmatism being sent by GCS. See Issue #1748. Despite our initial suspicions, this spurious astigmatism is not associated with the offsetting activities underway at the time. It seems to be a problem in the GCS WFS analysis since the problem is not seen with reanalysis by IDL.

Bad Seeing

UT06:00 The seeing has remained bad, so we are unable to to take the aberration measurements. Lucifer team is pursuing their tests.

UT07:17 GCS fails to identify the star in the acquisition image. See Issue #1749.

UT09:30 conditions seem to be improving.

BS9107
  • PA=0 luci_20081017_0127, 0128, and 129
  • PA=30 luci_20081017_130 and 131
  • PA=60 luci_20081017_132 and 133

AGW Orientation on-axis X=0.088 y=612.161 offsetPointing 120" south

UT10:55 Seeing improving, checking aberrations again
BS9107_GS33 with GS33 on axis
  • PA=93.5 luci_20081017_0135
  • PA=63.5 luci_20081017_0136
  • PA=33.5 luci_20081017_0137
  • PA=-26.5 luci_20081017_0138
  • PA=93.5 luci_20081017_0139

Seeing good again, Nancy reports astigmatism, redoing Andrew's tests
  • PA=86.0 luci_20081017_0144
  • PA=116.0 luci_20081017_0145
  • PA=146.0 luci_20081017_0146

11:45 More AGW orientation.
  • Procedure: after collim., stop WFS, pause guiding, offset AGW Y stage, offset IE, resume guiding.
  • Resume guiding failed three times in a row!

We did an interim test where we paused and resumed guiding and this worked. Ended up a start guiding worked for our purposes because it does not move the AGw stages.

Take astrometry data on LUCIFER for orientation UT12:19

  • Use LUCIastrom.cat:
    • BS9107 @ PA=270, GS07 (RA~03:32) - luci_20081017_0152 and 0153
    • Low-Galactic Latitude field (06:30 -20:00) at EL~36 luci_20081017_0154 and 0155

-- DavidThompson - 09 Oct 2008
Topic revision: r7 - 19 Oct 2008, JohnHill
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback